Ex 8.1 Class 7 Maths Question 1.
Find the ratio of:
(a) ₹ 5 to 50 paise
(b) 15 kg to 210 g
(c) 9 m to 27 cm
(d) 30 days to 36 hours

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(a) ₹ 5 to 50 paise Converting the given quantities into same units, we have ₹ 5 = 5 × 100 = 500 paise ∴ ₹ 5 : 50 paise = 500 paise : 50 paise [∵ ₹ 1 = 100 paise] = 10 : 1 So, required ratio is 10 : 1. (b) 15 kg to 210 g Converting the given quantities into same units, we have 15 kg = 15 × 1000 = 15000 g [∵ 1 kg = 1000 g] ∴ 15 kg : 210 g = 15000 g : 210 g = 1500 : 21 = 500 : 7 So, the required ratio is 500 : 7. (c) 9 m to 27 cm Converting the given quantities into same units, we have 9 m = 9 × 100 = 900 cm ∴ 9m: 27 cm = 900 cm : 27 cm [∵ 1 m = 100 cm] = 100 : 3 So, the required ratio is 100 : 3. (d) 30 days to 36 hours Converting the given quantities into same units, we have 30 days = 30 × 24 hours [ ∵ 1 day = 24 hours] = 720 hours ∴ 30 days : 36 hours = 720 hours : 36 hours = 20:1 So, the required ratio is 20 : 1.